Abstract Submitted to the XIV International Conference for Physics Students

Title: Elastic Peak Monitoring for the CLAS

Authors: DAVID C. VERMETTE, GERARD P. GILFOYLE, Physics Department, University of Richmond, VA, USA

Category: Nuclear Physics

We have developed code to monitor one of the large, particle detectors (CLAS) at the Thomas Jefferson National Accelerator Facility (TJNAF). TJNAF is a large electron accelerator designed to elucidate the quark nature of matter. The software collects data generated from collision events in the spectrometer and periodically analyzes it. The elastic scattering peak in the W spectrum is fitted to a Gaussian curve to measure its position and width. The position of the elastic peak is a measure of the accuracy of the detector calibration. The width of this peak is proportional to the resolution of the data. The fit procedure is subjected to several tests to ensure quality. A minimum number of counts is required in the elastic peak region and the range of the values of the uncertainties is restricted. These results are stored in a database and can be viewed over the internet. We will present plots of the elastic peak data and related parameters to analyze how the detector’s response can be monitored over time.